2.85 moles of an ideal gas with CV,m=3R/2 undergoes the transformations described in the following list from an initial state described by T = 310. K and P = 1.00 bar.

Part A:The gas is heated to 600 K at a constant volume corresponding to the initial volume. Calculate q for this process. Express your answer with the appropriate units.

Part B:The gas is heated to 600 K at a constant volume corresponding to the initial volume. Calculate w for this process. Express your answer with the appropriate units

Part C:The gas is heated to 600 K at a constant volume corresponding to the initial volume. Calculate ΔU for this process. Express your answer with the appropriate units.

Part D:The gas is heated to 600 K at a constant volume corresponding to the initial volume. Calculate ΔH for this process. Express your answer with the appropriate units.

Part E:The gas is heated to 600 K at a constant volume corresponding to the initial volume. Calculate ΔS for this process. Express your answer with the appropriate units.
Answer #1

Part A

The question asks to calculate for 'q' only

as there is no change in volume, work done = 0

so,

q = nCV

q = 2.85 * 3 * 8.314 / 2 * ( 600 - 310)

q = 10307.28 J

---------------------------------------------------------------------------

Part B

W = 0 because volume is constant

-----------------------------------------------------------------------

Part C

U = q

U =  10307.28 J

------------------------------------------------------------------------

Part D

H = 2.85 * 5 * 8.314 / 2 * ( 600 - 310)

H = 17178.8 J

--------------------------------------------------------------------------

Part E

S = nCv ln (T2 / T1)

S = 2.85 * 3 * 8.314 / 2 * ln ( 600 / 310)

S = 23.47 J/K